EDITORIAL

Terry Goodman

Recently, I had a discussion about mathematics with a number of adults — not a math major in the group. The discussion focused on the relative difficulty in learning mathematics compared to learning other subjects. Most of the group argued that mathematics was, by far, the most difficult and least meaningful subject that they had studied. This conclusion was voiced with not a little emotion and vigor.

Comments were made such as, “I never understood math.” “I just tried to memorize the rules and hope I could work the problems.” “Math just never made sense to me and I’m still afraid of it.” “I know math is supposed to be important, but I didn’t see how it was relevant to me.” “Only the ‘brains’ can learn math.” I must admit that this was not the first time I had heard such

comments. The adults in this particular group were college graduates (several with graduate degrees), successful, and bright. As I listened to their discussion, I tried to formulate some conjectures as to how they had come to have such negative feelings toward mathematics. Some may argue that mathematics is naturally difficult for people in general. In my experience with young children, however, I have found them to be eager and very capable mathematics learners.

I do have my theory (probably more of an opinion) as to why so many adults have a rather negative view of mathematics. Rather than subject you to this theory/opinion, I want to encourage you, our readers, to send us your thoughts on this subject. This may take the form of letters to the editor, guest editorials, or articles. I believe this sharing of information/ideas could be quite interesting and useful for all of us. We look forward to your comments.